Hello, I have a few equations that I can not solve. Can someon PLEASE help.
Thank you!

1. 4x-12
-----------
6x^2-18x

2. 6x^2+17x+5
---------------
3x^2+13X+4


3. Y^2 TIMES 4Y
---------------------
12Y TIMES Y-7


4. X^2 TIMES X^2-5X+6
----------------------
X^2-4 TIMES 3X-X^2


Answer by tutorcecilia(2144)

1. 4x-12
-----------
6x^2-18x

.
.
Factor:
4(x-3)/6x(x-3)
.
Cancel out wherever possible:
Answer: 2/3x
.

.
2. 6x^2+17x+5
---------------
3x^2+13X+4
.
.
Factor:
(3x+1)(2x+5)/(3x+1)(x+4)
.
.
Cancel out wherever possible:
(2x+5)/(x+4)
.
.

3. Y^2 TIMES 4Y
---------------------
12Y TIMES Y-7
.
Combine terms:
4y^3/12y(y-7)
.
Cancel out:
y^2/3(y-7)

.
.
.
4. X^2 TIMES X^2-5X+6
----------------------
X^2-4 TIMES 3X-X^2
.
Factor:
x^2(x-3)(x-2)over(x+2)(x-2)(x)(3-x)
.
Cancel out wherever possible:
(x)(x-3)/(x+2)(3-x)
.
Multiply the numerator and denominator by (-1) because: (3-x) = (-1)(3-x)=(-3+x)= (x-3)
(-1)(x)(x-3)/(-1)(x+2)(3-x)
.
.
(-1)(x)(x-3)/(-x-2)(x-3)
Cancel out:
(-1)x/(-x-2)
.
.
Factor out (-1):
(-1)(x)/(-1)x+2)
.
Cancel out:
x/x+2